# Dual-Model Weight Selection and Self-Knowledge Distillation for Medical Image Classification

### v1

@@ 6 Conclusion
Herein, we proposed a novel medical image classification method that combines dual-model weight selection with SKD. This integration enables small models to achieve high classification accuracy while maintaining low computational costs. Experimental results on publicly available datasets—chest X-ray, lung CT scan, and brain MRI datasets—demonstrate the effectiveness and robustness of our approach. Our method shows strong potential for deployment in real-world medical settings where computational resources are often limited. By enhancing the learning capacity of compact models without increasing memory or processing demands, it offers a practical and scalable solution for clinical applications. Future work should investigate its generalizability to other imaging domains, such as ultrasound or histopathology, to fully assess its applicability across a wide range of medical tasks.

### v2

@@ 6 Conclusion
In this work, we introduced a dual-model weight selection framework combined with self-knowledge distillation for medical image classification. By constructing complementary parameter subsets and employing an EMA-based auxiliary teacher, the proposed method enhances the representation capability of compact models while maintaining low computational cost. Experiments across three imaging modalities, i.e., chest X-ray, lung CT, and brain MRI, demonstrated consistent improvements in classification performance and stability over baseline approaches, highlighting the effectiveness and practicality of the framework in resource-constrained clinical scenarios. While the current evaluation focuses on publicly available datasets, further investigation is needed to assess the robustness of models under domain shifts, such as variations in scanner types and acquisition protocols. In future work, we aim to extend the method to additional medical imaging tasks (e.g., ultrasound and histopathology), explore adaptive strategies for dynamic subset construction, and perform cross-center validation to evaluate generalizability more comprehensively.